Traditional Product Costing Versus Activity-Based Costing
High Country Outfitters, Inc., makes backpacks for large sporting goods chains that are sold under the customers' store brand names. The accounting department has identified the following overhead costs and cost drivers for next year:

Overhead Item Expected Costs

Cost Driver

Maximum Quantity
Setup costs

$1,216,800

Number of setups

7,200
Ordering costs

312,000

Number of orders

60,000
Maintenance

1,200,000

Number of machine hours 80,000
Power

120,000

Number of kilowatt hours

600,000

Total predicted direct labor hours for next year is 60,000. The following data are for two recently completed jobs:

Job 201 Job 202
Cost of direct materials $15,000 $16,500
Cost of direct labor $20,625 $72,750
Number of units completed 1,125 915
Number of direct labor hours 270 330
Number of setups 18 22
Number of orders 24 45
Number of machine hours 540 450
Number of kilowatt hours 270 360

a. Determine the unit cost for each job using a traditional company-wide overhead rate based on direct labor hours.
Round rate to two decimal places.
Overhead rate per direct labor hour $Answer

Mark 0.00 out of 1.00

Use rounded overhead rate calculated above for calculations below. Round cost answers to the nearest whole number, when needed. Round cost per unit to two decimal places, if needed.

Job 201 Job 202
Direct materials Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Direct labor Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Overhead Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Total job cost Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Units Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Per unit cost Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

b. Determine the unit cost for each job using ABC. Activity-based overhead rates:

Round answers to two decimal places, if needed.

Setup costs

Answer

Mark 0.00 out of 1.00

per setup
Ordering costs

Answer

Mark 0.00 out of 1.00

per order
Maintenance costs Answer

Mark 0.00 out of 1.00

per machine hour
Power

Answer

Mark 0.00 out of 1.00

per kwh


NOTE: Round all answers (except cost per unit) to the nearest whole number, when applicable. Round cost per unit answers to two decimal places.  

Job 201 Job 202
Direct materials Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Direct labor Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Overhead
Setup cost Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Ordering costs Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Maintenance costs Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Power Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Total job costs Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Units produced Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00

Cost per unit Answer

Mark 0.00 out of 1.00

Answer

Mark 0.00 out of 1.00
Answer #1

(a) Unit cost for each job using a traditional company-wide overhead rate based on direct labor hours.

Overhead recovery rate= 2,848,800/60000= 47.48 per DL hour

Job 201

Job 202

Direct materials

          15,000

       16,500

Direct labor

          20,625

       72,750

Overhead

          12,820

       15,668

Total job cost

          48,445

    1,04,918

Units

            1,125

            915

Per unit cost

            43.06

       114.66

(B) Activity-based overhead rates

Setup costs

1,216,800/7200= 169 per setup

Ordering costs

312,000/60,000= 5.2 per order

Maintenance costs

1,200,000/80,000= 15 per machine hour

Power

120,000/600,000= 0.2 per kilowatt hour

Unit cost for each job under ABC

Job 201

Job 202

Direct materials

          15,000

       16,500

Direct labor

          20,625

       72,750

Overhead

Setup cost

3042

3718

Ordering costs

124.8

234

Maintenance costs

8100

6750

Power

54

72

Total job costs

          46,946

    1,00,024

Units produced

            1,125

            915

Cost per unit

            41.73

       109.32